OpenStudy (anonymous):

After the softball game, Alise's coach bought each of the 14 players a double-scoop or triple-scoop ice cream cone at the local ice cream shop. The coach bought a total of 37 scoops of ice cream. Use a model to determine how many girls got a triple-scoop ice cream cone. (A) 6 (B) 7 (C) 8 (D) 9

OpenStudy (anonymous):

D + T = 14 2D + 3T = 37

OpenStudy (anonymous):

Use a model to determine how many girls got a triple-scoop ice cream cone. Let p be the # of players that got the triple scoop37 = 3p + 2(14-p)37 = 3p + 28 – 2p37 – 28 = 3p – 2p9 = 1pThe answer is 9 girls got a triple-scoop, while the other 5 girls got double-scoop. Verify the number of scoops: 9(3) + 5(2) = 27 + 10 = 376.